OpenStudy (vannah):

what is the slope of the line containing the points (3,4) and (-6,10)?

OpenStudy (anonymous):

4-10 / 3 -- 6 -6/9 -2/3 that's slope y-y1=m(x-x1) y-4=-2/3(x-3)

OpenStudy (anonymous):

The equation for a slope is \[slope = \Delta y/ \Delta x\] so you do 10-4/-6-3 = 6/-9 = -2/3